OpenStudy (anonymous):

Subtract. (x + 1) - (-2x - 5)

OpenStudy (callisto):

(x + 1) - (-2x - 5) = (x + 1) - (-2x) - (- 5) = x +1 +2x +5 = x+2x +1 +5 =?

OpenStudy (anonymous):

it just requires you to simplify it.so 3x+6 is enough